Amanda is expected to receive full social security retirement benefit of $15500 annually when she turns 65. These benefits can be increased by 40% if she delays taking them till age 70. If her life expectancy is 85 years and expected rate of return is 3%, should Amanda take full benefit at 65 or delay benefit till 70? Assume no other income and taxes during retirement.

She should start at 65 since present value of benefits is $259,053 as against $223,462 at 70

She should wait till 70 since present value of benefits is $259,053 as against $230,601 at 65

She should wait till 70 since present value of benefits is $259,053 as against $223,462 at 65

She should start at 65 since present value of benefits is $230,601 as against $223,462 at 70
Answer #1

Present value of benefits if taken at 65=15500/3%*(1-1/1.03^20)=230601

Present value of benefits if delayed till 70=15500*1.4/3%*(1-1/1.03^15)*1/1.03^5=223462

She should start at 65 since present value of benefits is $230,601 as against $223,462 at 70